# Electronic structure of a nodal line semimetal candidate

### v1

@@ III Results and discussion
Next, we move onto the HS direction. Fig. 3(d), presents the dispersion cut along , its second derivative and its theoretically calculated surface spectrum (Figs. 3(e) and (f)).
@@
An electron like pocket is present along this direction (dispersion map along measured at other photon energies are presented in SM, which is characteristic of SbTe type materials [31, 38, 42]. Bulk band calculations and surface projected spectrum, both exhibit similar pocket like feature suggesting presence of surface and bulk bands forming the pocket.
@@
Proceeding to the dispersion cut along the , presented in Fig. LABEL:fig_04(a), two linear bands with different velocities converge to form an intersection, which is well reproduced in surface projected calculations presented at Fig. 4(c). This intersection near 200 meV, is a part of the nodal line which spans a whole plane surrounding the point, this loop like feature is fairly evident from the CECs presented in Fig. 2 as well. Again, contrasting the experimental observations, bulk calculation (Fig. 1(b)) suggests that, a gap occurs at the intersection of bands CB2 and VB2 (see Fig. 1(b)), which persists in a reduced form, if SOC is taken into account. This gap was also beyond detection limits of the ARPES detector. Another notable aspect is a hole pocket like feature which can also be seen centering the (or ) point in cuts parallel to both and directions (see Figs. 3 and 4 respectively), throughout all the incident photon energies (presented in SM), but its spectral intensity was highly sensitive to the incident photon energy.
@@
Thermodynamic, magnetic and transport measurements on TbSbTe single crystals coherently exhibit the AFM to PM phase transition near 5.1 K. The electronic band structure of TbSbTe, in ARPES spectrographs, exhibit two isocentric diamond shaped sheets around the point similar to other SbTe and ZrSiS type materials. For TbSbTe the double-sheet is consistently present throughout the FS, unlike heavier members of SbTe (for an example HoSbTe [59]), where the bands tend to gap out while radially going towards the direction. Another feature of this FS is, the bands are not as well separated as in NdSbTe [38] or PrSbTe [60, 42], nor coalesced into a single sheet like GdSbTe [31].
@@
The nodal line features observed in the cuts parallel to the are well reproduced in bulk electronic calculations. In space group , symmetries in combination with nonsymmorphic symmetry plays a crucial role in realizing enforced degeneracy in the vicinity of Fermi energy. Dirac like crossings observed in the HS directions parallel to
@@
and (which are not located at exact HS points X or M) are protected by symmetry, although consideration of SOC destabilizes these nodal crossings. Despite the theoretical predictions of destabilized nodal line features along and , in experimental observations, the nodal lines remain ungapped. This phenomenon could potentially be ascribed to constrained resolution within the vacuum ultraviolet region, resulting in the energy cuts encompassing a spectrum of values rather than being confined to a specific value. An alternate explanation might stem from the prevalent incongruity between experimental results and theoretical predictions often observed in 4 systems. On the other hand, nonsymmorphic (glide) symmetry protects the out of plane crossing along direction (along the HS direction)(See Fig. 1(b)), making TbSbTe a genuine nodal line semimetal candidate.
IV Conclusion
@@
To summarize, we present comprehensive temperature-dependent thermodynamic and electric transport investigations, which consistently indicates an AFM to PM transition at TN = 5.1 K. Concurrently, we studied the electronic band structure of a new member of the SbTe family, TbSbTe, using high-resolution ARPES combined with theoretical first-principles calculations. Our electronic structure study revealed multiple band intersections, including a Dirac like crossing along the HS direction, which is part of a nodal line extending along bulk direction. This crossing remains ungapped over , as established in both our theoretical work considering the effect of SOC and our experimental ARPES-based observations. These results implicate TbSbTe as a promising candidate for examining the interaction of symmetry, topology and SOC in SbTe series of materials.
Acknowledgements
@@
M. N. acknowledges support from the National Science Foundation under the CAREER award DMR-1847962 and the Air Force Office of Scientific Research for MURI Grant No. FA9550-20-1-0322. D. K. and T.R. were supported by the National Science Centre (Poland) under research grant 2021/41/B/ST3/01141. V. B. and S. R. acknowledge the support from Idaho National Laboratory’s Laboratory Directed Research and Development (LDRD) program under the Department of Energy (DOE) Idaho Operations Office Contract DEAC07-05ID14517. K. G. acknowledges support from the Division of Materials Science and Engineering, Office of Basic Energy Sciences, Office of Science of the U. S. Department of Energy (U.S. DOE). A. P. acknowledges the support by National Science Centre (NCN, Poland) under Projects No. 2021/43/B/ST3/02166 and also appreciates the funding in the frame of scholarships of the Minister of Science and Higher Education (Poland) for outstanding young scientists (2019 edition, No. 818/STYP/14/2019). This research utilized resources of the ALS at the Lawrence Berkeley National Laboratory, which is a DOE Office of Science User Facility under Contract No. DE-AC02-05CH11231. Pressure-dependent magnetic measurements were performed at the State University of New York (SUNY), Buffalo State, and supported by the National Science Foundation, Launching Early-Career Academic Pathways in the Mathematical and Physical Sciences (LEAPS-MPS) program under Award No. DMR-2213412. We express our gratitude to Dr. Sung-Kwan Mo for providing valuable support with the beamline at the Advanced Light Source (ALS).

### v2

@@ III Results and Discussion
Next, we move onto the -- HS direction. Fig. 3(d), presents the dispersion cut along --, its second derivative and its theoretically calculated surface spectrum (Figs. 3(e) and (f)).
@@
An electron like pocket is present along this direction (dispersion map along -- measured at other photon energies are presented in the SM Figs. S7(k-o) [66]), which is characteristic of SbTe type materials [33, 42, 46]. Bulk band calculations and surface projected spectrum, both exhibit similar pocket like feature suggesting presence of surface and bulk bands forming the pocket.
@@
Proceeding to the dispersion cut along the --, presented in Fig. 4(a), two linear bands with different velocities converge to form an intersection, which is well reproduced in surface projected calculations presented in Fig. 4(c). Again, contrasting the experimental observations, bulk calculation (Fig. 1(b)) suggests that, a gap occurs at the intersection of bands CB2 and VB2 (see Fig. 1(b)), which persists in a reduced form, if SOC is taken into account. This gap was also beyond the detection limits of the ARPES detector. Another notable aspect is a hole pocket like feature which can also be seen centering the (or ) point in cuts parallel to both and directions (see Figs. 3 and 4 respectively), throughout all the incident photon energies (presented in SM Fig. S7(a-e) and (f-j) respectively [66]), but its spectral intensity was highly sensitive to the incident photon energy.
Momentum dependent surface states parallel to -.
A detailed analysis of the surface states based on theoretical and experimental results are presented in Fig. 5. The theoretically calculated surface Green function and slab band structures are presented in Fig. panels 5(a-e), along and parallel to the - HS direction (see the cuts in the Fermi surface map in Fig. 5(f)). The grey bands corresponds to bulk and blue bands corresponds to surface originated bands.
The band features centering the and X HS points are predominately bulk originated. The theoretical calculations are well reproduced in the experimental ARPES based observations presented in Figs. 5(g-l) and their corresponding second derivatives in Figs. 5(m-r). Figs. 5(g) and h are ARPES dispersion cuts measured with 55 and 50 eV of incident photon energies along -, exhibiting surface bands forming pockets around the HS point. From to 600 meV binding energy, a linearly dispersing surface band is observed within the bulk bands. Below that, in the proximity of 800 meV binding energy, two branches of surface bands are observed. Another set of surface bands are observed right at the HS point. In total, we see four regions containing the surface bands. Cuts parallel to - directions, presented in Figs. 5(h-l) and their corresponding second derivative plots in 5(n-r), exhibit strong momentum dependence of the surface bands.
IV Discussion
@@
Thermodynamic, magnetic and transport measurements on TbSbTe single crystals coherently exhibit the AFM to PM phase transition near 5.8 K. The electronic band structure of TbSbTe, in ARPES spectrographs, exhibit two isocentric diamond shaped sheets around the point similar to other SbTe and ZrSiS type materials. For TbSbTe the double-sheet is consistently present throughout the FS, unlike heavier members of SbTe (for an example, HoSbTe [35]), where the bands tend to gap out while radially going towards the - direction. Another feature of this FS is that, the bands are not as well separated as in NdSbTe [42] or PrSbTe [68, 46], nor coalesced into a single sheet like GdSbTe [33].
@@
The features of the nodal line observed in the cuts parallel to the - are well reproduced in bulk electronic calculations. In space group , symmetries in combination with nonsymmorphic symmetry plays a crucial role in realizing enforced degeneracy in the vicinity of Fermi energy. This space group hosts a glide plane , screw axes, and , mirror symmetries , and , and the spatial inversion symmetry .
@@
Dirac crossings observed in the HS directions parallel to -X (with different incident photon energies, i.e. at different locations), right at the points, are degeneracies enforced by the combined effect of glide plane, and time reversal symmetry, . These crossings are part of the nodal line that extends along the X-R HS direction. The combination of these symmetries ensures the robustness of this nodal line even under consideration of the effect of SOC. The screw axis symmetries, in conjugation with the inversion symmetry, allows the other crossing (at the relatively higher energies), which also extends longitudinally along the axis. Mirror symmetry, allows the crossings in the -M HS direction, which are parts of the nodal line extending parallel to the M-A direction, often observed in the ZrSiS type systems. However, nodes in this direction were not observed conclusively in this study. Although, except for the nonsymmorphic symmetry enforced nodal line (along X-R, at the X point), consideration of SOC destabilizes the other nodal crossings. Despite the theoretical predictions of destabilized nodal line features, in experimental observations, the nodal lines remain ungapped. This phenomenon could potentially be ascribed to constrained resolution within the vacuum ultraviolet region, resulting in the energy cuts encompassing a spectrum of values rather than being confined to a specific value. An alternate explanation might stem from the prevalent incongruity between experimental results and theoretical predictions often observed in 4 systems. Moreover, while the previously discussed symmetries preserve the degeneracies in the PM phase, consideration of magnetic ordering (i.e. AFM or FM magnetic ordering) usually leads to breaking of these degeneracies. In the electrical and thermodynamic measurements, TbSbTe undergoes a PM to AFM phase transition, thus we restrict our considerations to AFM ordering. AFM spin orientation of Tb along (001) direction elongates the lattice parameter into (technically space group with ). This still guarantees the nodal line along the X-R direction. A similar analysis was also performed in Ref. [30].
V Conclusion
@@
To summarize, we present temperature and magnetic field-dependent thermodynamic and electric transport investigations, which consistently indicates an AFM to PM transition at K. Concurrently, we studied the electronic band structure of a new member of the SbTe family, TbSbTe, using high-resolution ARPES combined with theoretical first-principles calculations. Our electronic structure study revealed multiple band intersections, including a Dirac crossing, formed by bulk bands, in the -X HS direction, which is part of a nodal line extending along X-R bulk direction. This crossing remains ungapped over , as established in both our theoretical work considering the effect of SOC and our experimental ARPES-based observations. Another Dirac like crossing in this direction, was gapped in the theoretical calculations. These results implicate TbSbTe as a promising candidate for examining the interaction of symmetry, topology and SOC in SbTe series of materials.
acknowledgments
@@
M. N. acknowledges support from the National Science Foundation under the CAREER award DMR-1847962 and the Air Force Office of Scientific Research for MURI Grant No. FA9550-20-1-0322. N. V. acknowledges the support from the MPS Alliances for Graduate Education and the Professoriate–Graduate Research Supplements (AGEP–GRS). D. K. and T.R. were supported by the National Science Centre (Poland) under research grant 2021/41/B/ST3/01141. V. B. and S. R. acknowledge the support from Idaho National Laboratory’s Laboratory Directed Research and Development (LDRD) program under the Department of Energy (DOE) Idaho Operations Office Contract DEAC07-05ID14517. K. G. acknowledges support from the Division of Materials Science and Engineering, Office of Basic Energy Sciences, Office of Science of the U. S. Department of Energy (U.S. DOE). A. P. acknowledges the support by National Science Centre (NCN, Poland) under Projects No. 2021/43/B/ST3/02166.
This research utilized resources of the ALS at the Lawrence Berkeley National Laboratory, which is a DOE Office of Science User Facility under Contract No. DE-AC02-05CH11231. Magnetic measurements were performed at the State University of New York (SUNY), Buffalo State, and supported by the National Science Foundation, Launching Early-Career Academic Pathways in the Mathematical and Physical Sciences (LEAPS-MPS
) program under Award No. DMR-2213412. We express our gratitude to Dr. Sung-Kwan Mo for providing valuable support with the beamline at the Advanced Light Source (ALS).
Supplemental Materials for:
Electronic Structure of a Nodal Line Semimetal Candidate TbSbTe
SUPPLEMENTARY ARTICLES
S1. Single crystal characterization and theoretical band structure calculations
Fig. S1(a) presents a photograph of single crystals of TbSbTe grown for the purpose of the present study. They were found by XRD to have the expected tetragonal crystal structure with the lattice parameters close to those reported in the literature [1, 2]. The natural phase of the as-grown crystals is (001), as proven by the Laue technique (see Fig. S1(b)). The XRD and EDX data indicated that the crystals are single-phase and homogeneous with the chemical composition very close to the ideal one (viz. S1(c)). Figs. S2(a-c) visualizes the crystallographic unit cell of TbSbTe that can be represented as a stacking of Tb-Te stabs and quadratic Sb layers. The bulk Brillouin zone, its projection on the (001) plane and the positions of the nodal lines (along with the symmetries protecting/enforcing them) are shown in Fig. S2(d). The nodal lines along the X–R high symmetry direction, are nonsymmorphic symmetry enforced crossings [glide plane symmetry, in combination with time reversal symmetry, ], which are represented by the green broken lines in Figs. S2(d). The allowed nodal line along X–R [screw axis, and inversion symmetry, ] is represented with yellow broken lines. The corresponding Dirac nodes are indicated with green and yellow dots in the (001) surface Brillouin zone [3]. To better understand the positions of the nodal lines, visualizations are presented in Figs. S2(e) and (f). The calculated bulk band structure of TbSbTe, in the nonmagnetic and antiferromagnetic (AFM) phase is presented in the Fig. S3(a) and (b) respectively. Plokhikh et al. reported the different magnetic structures of TbSbTe in details in Ref. [4]. Such magnetic configurations are exclusively stabilized in large unit cells, where magnetic moments arrange into square blocks exhibiting AFM-like ordering. However, in the present case, these magnetic structures proved to be numerically unstable. The results presented in Fig. S3(b) are based on DFT+U, where the effective correlation parameter (U), shifts the bands away from the Fermi energy, appearing as flat bands ranging from -2 eV to -4 eV and 0.5 eV to 2 eV. Thus, the magnetic schemes have minimal effect on the low lying electronic structure of the material. One significant change that occurs to the enforced band crossing at the X point at the AFM phase with SOC considered, is the crossing is gapped. The nodal lines obtained from different band crossings are visualized in Fig. S4.
S2. ARPES based band structure measurements
Angle resolved photoemission spectroscopy (ARPES) based measurements were carried out at Advanced Light Source (ALS) beamline endstation 10.0.1.1 at 18 K. The sample crystals were cleaved at ultra high vacuum better than Torr. For the measurements of the ARPES maps, the energy resolution was set to be better than 20 meV and the angle resolution was better than 0.2°. The constant energy contours ranging from meV to meV, measured with 60 eV incident photon energy, are presented in Fig. S5, which is a continuation of Fig. 2 of the main manuscript. The Fermi surfaces and the constant energy contours measured using 45 eV to 55 eV incident photon energies are displayed in SM Fig. S6.
Fermi surface at all the photon energies persistently exhibit the double-sheet nature similar to the 60 eV data presented in the main manuscript. Fig. S7 presents a detailed analysis of ARPES data to show the dual sheet nature of the Fermi surface. Intensity profile was taken from two energy positions from the cut parallel to the direction (Fig. S7(a)). The corresponding Fermi surface maps (cut C1) at two different energies (with h = 55 and 60 eV) are presented in Fig. S7(b) and CEC corresponding to cut C2 is presented in Fig. S7(c). Intensity profile taken from the broken red line in Fig. S7(b), shown in Fig. S7(d), exhibits two peaks, representing the two sheets, discussed above.
Electronic dispersion maps measured with incident photon energies ranging from 45 eV to 65 eV are presented in SM Fig. S8. SM Figs. S8(a-e) show crossings in the dispersion cut parallel to - direction which are part of nodal lines along direction at approximately meV and meV.
In SM Figs. S8(f-j), the crossing near meV is persistent along increasing incident photon energy along . All of the aforementioned cuts host a strongly energy dependent hole pocket centered at . In the energy range from the Fermi level down to eV, SM Fig. S8(k–o) reveals a pronounced energy-dependent Dirac node feature along the – momentum cuts. The variation in intensity with incident photon energy indicates that these features arise from bulk electronic states. In contrast, the wing-like feature observed below eV remains invariant with respect to photon energy, suggesting a surface-derived origin.
SM Fig. S9 presents the ARPES dispersion map and second derivative of the dispersion cut at 50 eV, near - direction. To better visualize the band features, the stacked momentum distribution curves (MDC) integrated within a span of 20 meV are presented in SM Fig. S9(c). Second derivative plots of the dispersion cuts at different incident photon energies are presented in SM Fig. S10, and the positions of the crossings are presented in SM Fig. S11, as a function of incident photon energy. The crossings which are not protected by the nonsymmorphic symmetry were gapped in the theoretical bulk calculations; magnified views of those crossings are presented in Fig. S12. The EDC taken at the corresponding position of the crossing (0.01 Å-1), shown in the Fig. S12(b) do not show any signature of gap opening, may be because of instrument resolution. The crossing appears to be slightly tilted, suggesting a type II nature. In Fig. S12(c), we calculated the MDCs of Fig. S12(a), and tracked the peak points connecting them (represented by blue broken lines).
Supplementary References
- [1] F. Gao, J. Huang, W. Ren, H. Wu, M. An, X. Wu, L. Zhang, T. Yang, A. Wang, Y. Chai, X. Zhao, T. Yang, B. Li, and Z. Zhang, Adv. Quantum Technol. 6, 2200163.
- [2] I. Plokhikh, V. Pomjakushin, D. J. Gawryluk, O. Zaharko, and E. Pomjakushina, Inorg. Chem. 61, 11399 (2022).
- [3] A. Topp, R. Queiroz, A. Grüneis, L. Müchler, A. W. Rost, A. Varykhalov, D. Marchenko, M. Krivenkov, F. Rodolakis, J. L. McChesney, B. V. Lotsch, L. M. Schoop, and C. R. Ast, Phys. Rev. X 7, 041073 (2017).
- [4] I. Plokhikh, V. Pomjakushin, D. J. Gawryluk, O. Zaharko, and E. Pomjakushina, J. Magn. Magn. Mater. 583, 171009 (2023).
SUPPLEMENTARY FIGURES